OpenStudy (anonymous):

Compute Δy and dy for x = 4 and dx = Δx = 1. y=5/x. I get dy=-1/4 and Δy=-5/16. What am i doing wrong??

OpenStudy (anonymous):

\[\Delta y=\frac{5}{4+1}-\frac{5}{4}=1-\frac{5}{4}=-\frac{1}{4}\] if i am not mistaken

OpenStudy (anonymous):

whereas \[\frac{dy}{dx}=-\frac{5}{x^2}\] \[dy=-\frac{5}{x^2}dx\] now replace x by 4 and get \[-\frac{5}{16}\] so i guess you just had them backwards